What is the equation of the line that is perpendicular to -x + y = 7 and passes through (-1, -1)?
a)-x + y = 2
b)-x – y = 2
c)x – y = 2
d)-x – y = 0
e)x + y = 0

Answer by rothauserc(4718)

-x + y = 7
standard slope - intercept form for this equation is
y = x + 7
slope is 1 and the slope of the line perpendicular is -1 (negative reciprocal)
so far we have
y = -x + b
now use the point (-1, -1)
-1 = -(-1) + b
-1 = 1 + b
b = -2
y = -x - 2
add x to both sides of =
x + y = -2
multiply both sides of = by -1
-x -y = 2
answer is b